# On the Selection Stability of Stability Selection and Its Applications

### v1

@@ Abstract
Stability selection is a widely adopted resampling-based framework for high-dimensional structure estimation and variable selection. However, the concept of ‘stability’ is often narrowly addressed, primarily through examining selection frequencies, or ‘stability paths’. This paper seeks to broaden the use of an established stability estimator to evaluate the overall stability of the stability selection framework, moving beyond single-variable analysis. We suggest that the stability estimator offers two advantages: it can serve as a reference to reflect the robustness of the outcomes obtained and help identify an optimal regularization value to improve stability. By determining this value, we aim to calibrate key stability selection parameters, namely, the decision threshold and the expected number of falsely selected variables, within established theoretical bounds. Furthermore, we explore a novel selection criterion based on this regularization value. With the asymptotic distribution of the stability estimator previously established, convergence to true stability is ensured, allowing us to observe stability trends over successive sub-samples. This approach sheds light on the required number of sub-samples addressing a notable gap in prior studies. The stabplot package is developed to facilitate the use of the plots featured in this manuscript, supporting their integration into further statistical analysis and research workflows.
Keywords: Bioinformatics, Feature Selection, Model Selection, Structure Estimation, Variable Selection
@@ 1 Introduction
The estimation of discrete structures, including graphs, clusters, or selection of variables, represents a long-standing and fundamental problem in statistics (Meinshausen and Bühlmann, 2010). Variable selection is a critical phase in the modeling pipeline and a fundamental step in data preprocessing, where raw data is refined to ensure its suitability for analysis or modeling. The objective is to eliminate variables that are noisy, redundant, or irrelevant (Tibshirani, 1996; Müller and Welsh, 2010), addressing a key challenge in contemporary data science (Nogueira et al., 2018).
@@
Stability selection, a widely recognized resampling-based variable selection framework, performs variable selection by examining how frequently variables are chosen across multiple randomly selected resamples (Meinshausen and Bühlmann, 2010). The selection frequency indicates how often a variable is included in the estimated models over various resamples. Meinshausen and Bühlmann (2010) analyzed the selection frequencies for individual variables over a range of regularization values, presenting the results in a plot known as the ‘stability path’. This plot facilitates the identification of variables that are consistently selected over the regularization grid.
In this paper, we shift the focus from examining the stability of individual variables across a grid of regularization values to evaluating the overall stability of the entire stability selection framework across the regularization grid. Employing an established stability estimator, we assess the stability of the framework in its entirety and introduce what we call ‘Stable Stability Selection’, which identifies the smallest regularization value that yields highly stable outcomes.
We illustrate the stable stability selection methodology through variable selection in the context of linear regression. However, the applicability of this approach extends beyond linear regression, as stability selection is not confined to this specific problem. Although detailed exploration is beyond the scope of this paper, stable stability selection is equally suitable for more complex regression-type models.
The stability of a variable selection method refers to the consistency with which it selects variables across different training sets drawn from the same underlying distribution (Kalousis et al., 2007). Evaluation of the stability of a statistical model by random resampling has been a well-established practice for decades. For example, Altman and Andersen (1989) applied this approach in Cox’s proportional hazards regression model (Cox, 1972) to identify stable variables, which were defined as those consistently selected with higher frequencies across multiple resamples.
@@
In the seminal work of Meinshausen and Bühlmann (2010), the stability selection framework was introduced, which offers a transformative method to identify stable variables in structure estimation tasks. The stability selection approach described by Meinshausen and Bühlmann (2010) uses sub-samples that are half the size of the sample size of the original dataset to carry out variable selection. Stability selection enables the identification of variables that are consistently recognized as important across the majority of sub-samples given a regularization value, considering them as stable variables. Furthermore, Meinshausen and Bühlmann (2010) established an asymptotic upper-bound for the Per-Family Error Rate, which represents the expected number of falsely selected variables. This bound relies on the assumptions of exchangeability in the distribution of falsely selected variables and that the selection algorithm performs no worse than random guessing.
Shah and Samworth (2013) introduced the complementary pairs stability selection method, which computes selection frequencies by counting how often a variable is included in models fitted to complementary $50\%$ sub-samples. Shah and Samworth (2013) introduced more rigorous upper-bounds for the expected number of variables selected with low selection probabilities.
It is somewhat surprising that, despite the framework being designated as ‘stability’ selection and the considerable attention stability selection has received since its inception, the overall stability of the framework itself has been largely overlooked, with no dedicated studies specifically addressing this critical aspect.
There is a substantial body of research on the stability of variable selection methods, where numerous stability measures have been proposed. For more comprehensive information, we refer to Kuncheva (2007), Nogueira et al. (2018), and Sen et al. (2021).
@@
Nogueira et al. (2018) conducted an extensive literature review on the topic, consolidating the desirable properties of stability measures into five mathematical conditions: fully defined, strict monotonicity, constant bounds, maximum stability occurs if and only if the selection is deterministic and correction for chance. A brief overview of these properties will be provided in the following paragraph. Nogueira et al. (2018) showed that none of the stability measures previously introduced satisfied all five conditions. In response, Nogueira et al. (2018) proposed a novel stability measure, which was the first in the literature to meet all these criteria, generalizing several previous works such as in Kuncheva (2007).
@@
Here, we give a concise review of the five conditions proposed by Nogueira et al. (2018). The first property stipulates that the stability measure should function effectively regardless of variations in the number of selected variables, meaning that it should not be limited to scenarios where the number of selected variables remains constant across resamples. The second property asserts that the stability measure should be a strictly decreasing function of the sample variance of the binary selection status of each variable across random resamples; as the variance in the selection status of each variable increases, the stability of the framework diminishes. The third property states that the stability measure should be bounded by specific constants, which facilitates the comparison of the stability of different procedures. The fourth property implies that maximum stability is achieved when the selection results across all resamples are identical, and vice versa. Finally, the fifth property posits that if all possible selection results are equally likely to occur across random resamples, the expected value of the stability measure should remain constant. This indicates that the stability measure should not be influenced by random patterns that may arise during the resampling process.
Nogueira et al. (2018) conceptualized the stability of a variable selection procedure as an underlying population parameter that requires estimation. In addition, Nogueira et al. (2018), linking their work with Gwet (2008), demonstrated that as the number of resamples approaches infinity, their stability measure, $\hat{\Phi}$, converges to a Normal distribution with a location parameter $\mu=\Phi$, where $\Phi$ denotes population stability. Consequently, $\hat{\Phi}$ is a consistent estimator of the population stability $\Phi$; that is, as the number of resamples approaches infinity, $\hat{\Phi}$ converges in probability to the true population stability $\Phi$. This property allows us to construct confidence intervals and perform hypothesis testing on the stability values.
@@
Nogueira et al. (2018) compared the stability of the stability selection with that of the Least Absolute Shrinkage and Selection Operator (LASSO; Tibshirani, 1996), concluding that the former exhibits greater stability than the latter in most of the scenarios considered.
@@
In this paper, our aim is to apply the stability estimator proposed by Nogueira et al. (2018) to evaluate the stability of stability selection. Furthermore, we proxify the stability values to find the optimal regularization value to achieve high stability with the least possible loss in accuracy, if applicable. The optimal regularization value can be employed to calibrate two critical parameters of stability selection: the decision-making threshold and the expected number of falsely selected variables, balancing them with respect to one another. We also use the optimal regularization value to define a new selection criterion, called stable stability selection, prioritizing the selection of variables associated with highly stable outcomes, rather than those mostly selected in their best-case scenarios over the regularization grid, which is the primary criterion in the traditional stability selection framework.
In addition, the convergence value of the stability estimator serves as a reference to assess the reliability of the results obtained. If the stability selection process demonstrates instability, it is unreasonable to expect that its selections and preferences will be reliable. The point at which convergence occurs provides valuable insight into the required number of sub-samples, an aspect for which we found no dedicated research addressing its determination.
@@
Nogueira et al. (2018) suggested that stability and accuracy can be interpreted through the lens of a Pareto front Pareto (1896). We refer to this methodology as stability-accuracy selection. Building on this perspective, we demonstrate that, given two reasonable assumptions, our proposed regularization value constitutes a Pareto-optimal solution within this context.
Alternative stability estimators, aside from that proposed by Nogueira et al. (2018), may also be applicable; however, a detailed comparison and evaluation of different stability measures in the context of stability selection falls outside the scope of this paper.
@@
The structure of this paper is organized as follows. Section 2 outlines the proposed methodology. Section 3 describes the real and synthetic datasets used in the paper. The results of the application of the method to real and synthetic data are presented in Section 4. Finally, Section 5 summarizes and concludes the paper.
@@ 2 Methodology
In this section, we present a methodology to evaluate the selection stability of the stability selection framework and to determine the optimal regularization value accordingly. Although we demonstrate the proposed method within the context of stability selection outlined by Meinshausen and Bühlmann (2010), it is important to note that the methodology is adaptable to other formulations of stability selection, such as those described by Shah and Samworth (2013) and Beinrucker et al. (2016). We will introduce our methodology following a brief overview of the foundational approach established by Meinshausen and Bühlmann (2010).
@@
Variable selection, in accordance with the terminology of Meinshausen and Bühlmann (2010), typically involves categorizing covariates into two distinct groups: the signal group $S=\{k\neq 0\mid\beta_{k}\neq 0\}$ and the noise group $N=\{k\neq 0\mid\beta_{k}=0\}$, where $S\cap N=\emptyset$. The primary objective of variable selection is to accurately identify the signal group $S$.
@@
where $\lambda\in\mathbb{R}^{+}$ represents the LASSO regularization parameter. The set of non-zero coefficients can be identified as $\hat{S}(\lambda)=\{k\neq 0\mid\hat{\beta}_{k}(\lambda)\neq 0\}$, derived from solving the LASSO equation through convex optimization.
Although alternative methods exist for variable selection in linear regression models, a detailed exploration of techniques for determining the regularization method and penalty term in stability selection lies beyond the scope of this paper.
@@
$\hat{S}^{\text{stable}}=\{j\mid\max_{\lambda\in\Lambda}(\hat{\Pi}_{j}^{\lambda})\geq\pi_{\text{thr}}\};\quad j=1,\ldots,p,$ (1)
where $\Lambda$ represents the set of regularization values, $10<\pi_{\text{thr}}<$ 1 is the threshold for decision-making in variable selection, and $\hat{\Pi}_{j}^{\lambda}$ denotes the selection frequency of the $j$th variable given the regularization parameter $\lambda$.
Equation 1 identifies variables whose selection frequencies exceed $\pi_{\text{thr}}$ under the regularization value that maximizes their selection frequencies; therefore, it considers best-case of each variable for decision-making.
@@
We now move forward to stable stability selection. To implement this, we first define a grid of regularization values, $\Lambda$, from which the optimal $\lambda$ for LASSO is to be identified. For this purpose, we use as the default choice the $\lambda$ values generated by the cv.glmnet function from the glmnet package in R (Friedman et al., 2010), applying a 10-fold cross-validation to the complete dataset $\mathcal{D}$. Alternative regularization values can also be employed. However, a comparison of different methods for generating the regularization grid falls outside the scope of this paper. The goal is to identify the optimal regularization value $\lambda_{\text{stable}}\in\Lambda$ that yields highly stable outcomes with the least possible loss in terms of predictive ability.
The following procedure is applied to each regularization value $\lambda\in\Lambda$. In each iteration of the stability selection process, a random sub-sample, comprising half the size of the original dataset $\mathcal{D}$, is randomly selected. The LASSO model is then fitted to this sub-sample using the fixed value of $\lambda$, and the binary selection outcomes are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the number of sub-samples. Therefore, in the end, we obtain $|\Lambda|$ distinct selection matrices $M(\lambda)$.
In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is identified as part of the signal set $\hat{S}(\lambda)$ when LASSO is applied to the $b$th sub-sample given $\lambda$. In contrast, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is classified as belonging to the noise set $\hat{N}(\lambda)$ after applying LASSO to the $b$th sub-sample given $\lambda$.
@@
$\hat{\Phi}(M(\lambda))=1-\frac{\frac{1}{p}\sum_{j=1}^{p}s_{j}^{2}}{\frac{q(\lambda)}{p}\left(1-\frac{q(\lambda)}{p}\right)};\quad\lambda\in\Lambda,$ (2)
where $s_{j}^{2}$ denotes the unbiased sample variance of the binary selection statuses of the $j$th variable, while $q(\lambda)$ denotes the average number of variables selected under the regularization parameter $\lambda$. The stability estimator $\hat{\Phi}$ is bounded by $\left[-\frac{1}{B-1},1\right]$ (Nogueira et al., 2018); the larger the $\hat{\Phi}$, the more stable is $M(\lambda)$.
To interpret the stability values of Equation (2), Nogueira et al. (2018) adopted the guidelines proposed by Fleiss et al. (2004). According to Nogueira et al. (2018), stability values that exceed $0.75$ indicate excellent agreement between selections beyond what would be expected by random chance, while values below $0.4$ signify poor agreement between them. Based on Nogueira et al. (2018), stability values that reside between these thresholds are categorized as indicative of intermediate to good stability.
@@
$\lambda_{\text{stable}}=\min\left\{\lambda\in\Lambda\mid\hat{\Phi}(M(\lambda))\geq 0.75\right\}.$ (3)
The threshold value of $0.75$ is somewhat arbitrary and may not be attainable in certain situations. We address this limitation in the following paragraph.
As we will see in Section 4, $\lambda_{\text{stable}}$ may not exist in certain practical applications due to procedural instability, which prevents the stability values from exceeding $0.75$. In such cases, we propose
$\lambda_{\text{stable-1sd}}=\min\left\{\lambda\in\Lambda\mid\hat{\Phi}(M(\lambda))\geq\max_{\lambda\in\Lambda}\hat{\Phi}(M(\lambda))-\text{SD}_{\lambda\in\Lambda}\left(\hat{\Phi}(M(\lambda))\right)\right\},$ (4)
where SD represents the standard deviation of the stability values.
Both upper-bounds of the Per-Family Error Rate, mentioned in Section 1, provided by Meinshausen and Bühlmann (2010) and Shah and Samworth (2013), depend on the values of $\lambda$ and $\pi_{\text{thr}}$. The calibration of these two values requires the arbitrary selection of one of these two parameters, which can be challenging to justify. By adopting the optimal regularization value, that is $\lambda=\lambda_{\text{stable}}$ or $\lambda=\lambda_{\text{stable-1sd}}$, the upper-bound can be explicitly tailored to the data owner’s preferences. If a pre-determined $\pi_{\text{thr}}$ is chosen, the Per-Family Error Rate is deterministically obtained; alternatively, if a fixed pre-determined Per-Family Error Rate value is desired, the corresponding $\pi_{\text{thr}}$ is enforced.
@@
where $q_{\Lambda}$ represents the average number of selected variables over the regularization grid $\Lambda$. In a manner consistent with the approach of Bodinier et al. (2023), we can derive a point-wise control version of Equation 5 by restricting $\Lambda$ to a single value, $\lambda$, without affecting the validity of the original equation. Consequently, we can replace $q_{\Lambda}$ with $q(\lambda)$, as introduced in Equation 2. By employing $\lambda_{\text{stable}}$ within this formula and utilizing $q(\lambda_{\text{stable}})$, we establish a two-way control: fixing $\pi_{\text{thr}}$ allows us to determine the upper-bound, and vice versa. A similar approach can be applied to the upper-bounds proposed by Shah and Samworth (2013).
@@
$\hat{S}^{\text{stable}}=\left\{j\mid\hat{\Pi}_{j}^{\lambda_{\text{stable}}}\geq\pi_{\text{thr}}\right\};\quad j=1,\ldots.p.$ (6)
Equation (6) represents the variables with selection frequencies higher than $\pi_{\text{thr}}$ under the $\lambda_{\text{stable}}$, which corresponds to the highly stable outcomes. We term this selection criterion stable stability selection because it leverages highly stable outcomes to facilitate variable selection. If $\lambda_{\text{stable}}$ does not exist, $\hat{S}^{\text{stable-1sd}}$ can be defined similarly.
As outlined in Section 1, as the number of sub-samples increases, the convergence of $\hat{\Phi}$ to the true stability $\Phi$ is ensured. To illustrate this, we define $M^{(t)}(\lambda)$ as the matrix containing the selection outcomes from the first $t$ sub-samples for a given $\lambda\in\Lambda$, that is, the first $t$ rows of $M(\lambda)$. The objective is to evaluate the stability of $M^{(t)}(\lambda)$ across successive sub-samples in order to determine the appropriate cut-off point for the process, that is, the number of sub-samples required to achieve convergence in stability. Beyond this threshold, additional sub-samples do not significantly change the stability of $M(\lambda)$. We propose plotting the stability values of the selection matrix over the sequential sub-sampling, that is $\hat{\Phi}(M^{(t)}(\lambda));\;t=2,3,\ldots,B$, against $t$ to monitor the convergence status of the stability values. Given that the asymptotic distribution of $\hat{\Phi}$ is established, a confidence bound can be drawn along the curve to reflect the inherent uncertainty of the stability estimator.

### v2

@@ Abstract
Stability selection is a widely adopted resampling-based framework for high-dimensional variable selection. This paper seeks to broaden the use of an established stability estimator to evaluate the overall stability of the stability selection results, moving beyond single-variable analysis. We suggest that the stability estimator offers two advantages: it can serve as a reference to reflect the robustness of the results obtained, and help identify an optimal regularization value to improve stability. By determining this value, we calibrate key stability selection parameters, namely, the decision threshold and the expected number of falsely selected variables, within established theoretical bounds. The asymptotic distribution of the stability estimator allows us to observe convergence of stability values over successive sub-samples. This approach sheds light on the required number of sub-samples addressing a notable gap in prior studies. Pareto optimality of the proposed regularization value is also discussed. The stabplot R package is developed to facilitate the use of the plots featured in this manuscript, supporting their integration into further statistical analysis and research workflows.
Keywords: Bioinformatics, Feature Selection, Hyperparameter Tuning, Lasso, Variable Selection
@@ 1 Introduction
Stability selection, a widely recognized resampling-based variable selection framework, performs selection by examining how frequently variables are chosen across multiple randomly selected resamples (Meinshausen and Bühlmann, 2010). Meinshausen and Bühlmann (2010) analyzed the selection frequencies for individual variables over a range of regularization values, presenting the results in a plot known as the ‘stability path’. This plot facilitates the identification of variables that are consistently selected over the regularization grid. In this paper, we shift the focus from examining the stability of individual variables to evaluating the overall stability of the stability selection results. Employing an established stability estimator, we assess the stability of the results in its entirety and introduce what we call ‘Stable Stability Selection’, which identifies the smallest regularization value that yields highly stable results.
@@
The stability of a variable selection method refers to the consistency with which it selects variables across different training sets drawn from the same underlying distribution (Kalousis et al., 2007). The stability selection approach described by Meinshausen and Bühlmann (2010) uses sub-samples that are half the size of the sample size of the original dataset to perform variable selection. Stability selection enables the identification of variables that are consistently recognized as important across the majority of sub-samples given a regularization value, considering them as stable variables. Furthermore, Meinshausen and Bühlmann (2010) established an asymptotic upper-bound for the Per-Family Error Rate, which represents the expected number of falsely selected variables. Later, Shah and Samworth (2013) introduced the complementary pairs stability selection method, which computes selection frequencies by counting how often a variable is included in models fitted to complementary $50\%$ sub-samples. Shah and Samworth (2013) introduced more rigorous upper-bounds for the expected number of variables selected with low selection frequencies.
@@
There is a substantial body of research on the stability of variable selection methods, where numerous stability measures have been proposed. For more comprehensive information, we refer to Kuncheva (2007), Nogueira et al. (2018), and Sen et al. (2021). Nogueira et al. (2018) conducted an extensive literature review on the topic, consolidating the desirable properties of stability measures into five mathematical conditions. A brief overview of these properties will be provided in the following paragraph. Nogueira et al. (2018) showed that none of the stability measures previously introduced satisfied all five conditions. In response, Nogueira et al. (2018) proposed a novel stability measure, which was the first in the literature to meet all these criteria, generalizing several previous works such as in Kuncheva (2007).
@@
Here, we give a concise review of the five conditions proposed by Nogueira et al. (2018). The first property stipulates that the stability measure should function effectively regardless of variations in the number of selected variables. The second property asserts that the stability measure should be a strictly decreasing function of the sample variance of the binary selection status of each variable across random resamples. The third property states that the stability measure should be bounded by specific constants. The fourth property implies that maximum stability is achieved when the selection results across all resamples are identical, and vice versa. Finally, the fifth property posits that if all possible selection results are equally likely to occur across random resamples, the expected value of the stability measure should remain constant.
@@
Nogueira et al. (2018), linking their work with Gwet (2008), demonstrated that as the number of resamples approaches infinity, their stability measure, $\hat{\Phi}(\cdot)$, converges to a Normal distribution with a location parameter $\mu=\Phi$, where $\Phi$ denotes population stability. They also compared the stability of the stability selection with that of the Least Absolute Shrinkage and Selection Operator (Lasso; Tibshirani, 1996), concluding that the former exhibits greater stability than the latter in most of the scenarios considered.
@@
In this paper, we apply the stability estimator proposed by Nogueira et al. (2018) to the stability selection results, and proxify the stability values to find the optimal regularization value that achieves high stability with the least possible loss in prediction accuracy, if applicable. The optimal regularization value then is employed to calibrate two critical parameters of stability selection: the decision-making threshold and the expected number of falsely selected variables, balancing them with respect to one another. We also use the optimal regularization value to define a new variable selection criterion, called stable stability selection, prioritizing the selection of variables associated with highly stable results, rather than those mostly selected in their best-case scenarios over the regularization grid, which is the primary criterion in the traditional stability selection framework. In addition, the convergence value of the stability estimator serves as a reference to assess the reliability of the results obtained. The point at which convergence occurs provides valuable insight into the required number of sub-samples, an aspect for which we found no dedicated research addressing its determination.
@@
Nogueira et al. (2018) suggested that stability and prediction accuracy can be interpreted through the lens of a Pareto front Pareto (1896). We refer to this methodology as stability-accuracy selection. Building on this perspective, we demonstrate that, given two justifiable assumptions, the proposed regularization value constitutes a Pareto-optimal solution within this context.
The use of stability to enhance variable selection has been a subject of interest since at least the work of Liu et al. (2010). Among the existing literature, the study by Sun et al. (2013) is perhaps the most closely related to the present work and has continued to attract interest, as demonstrated by contributions such as Yang et al. (2020), Reavie (2021), Wen et al. (2023), and Lee and Chen (2024). In Sun et al. (2013), the regularization parameter of penalized regression models is selected based on a stability-driven criterion. Nonetheless, our methodology differs from theirs in several key aspects. Most notably, the stability measure they employ, originally introduced by Cohen (1960), is based solely on the intersection of some sets and does not account for selection variance. As argued by Nogueira et al. (2018), a stability measure should decrease as selection variance increases, a property not satisfied by the measure used in Sun et al. (2013). Furthermore, our approach is embedded directly within the stability selection framework, whereas the method in Sun et al. (2013) is designed to be used externally. Another contribution of this paper, which distinguishes it from Sun et al. (2013), is the use of stability values to inform the required number of sub-samples within the stability selection framework.
@@
The structure of this paper is organized as follows. Section 2 outlines the proposed methodology. Section 3 describes the real and synthetic datasets used in the paper. The results of applying the proposed method, along with the demonstration of Pareto optimality of the proposed regularization value, are presented in Section 4.
@@ 2 Methodology
In this section, we present a methodology to evaluate the selection stability of the stability selection results and to determine the optimal regularization value accordingly. Although we demonstrate the proposed method within the context of stability selection outlined by Meinshausen and Bühlmann (2010), it is important to note that the methodology is adaptable to other formulations of stability selection, such as those described by Shah and Samworth (2013). We will introduce our methodology following a brief overview of the foundational approach established by Meinshausen and Bühlmann (2010).
@@
Variable selection, in accordance with the terminology of Meinshausen and Bühlmann (2010), typically involves categorizing covariates into two distinct groups: the signal group $S:=\{k\neq 0\mid\beta_{k}\neq 0\}$ and the noise group $N:=\{k\neq 0\mid\beta_{k}=0\}$, where $S\cap N=\emptyset$. The primary objective of variable selection is to accurately identify the signal group $S$.
@@
where $\lambda\in\mathbb{R}^{+}$ represents the Lasso regularization parameter. The set of non-zero coefficients can be identified as $\hat{S}(\lambda):=\{k\neq 0\mid\hat{\beta}_{k}(\lambda)\neq 0\}$, derived from solving the Lasso equation through convex optimization.
@@
$\hat{S}^{\text{stable}}:=\{j\mid\max_{\lambda\in\Lambda}(\hat{\Pi}_{j}^{\lambda})\geq\pi_{\text{thr}}\};\quad j=1,\ldots p,$ (1)
where $\Lambda$ represents the set of regularization values, $\pi_{\text{thr}}\in[0.6,0.9]$ is the threshold for decision-making in variable selection, and $\hat{\Pi}_{j}^{\lambda}$ denotes the selection frequency of the $j$th variable given the regularization parameter $\lambda$. Equation (1) identifies variables whose selection frequencies exceed $\pi_{\text{thr}}$ under the regularization value that maximizes their selection frequencies; therefore, it considers best-case of each variable for decision-making. This approach has been criticized in subsequent studies, notably by Zhou et al. (2013).
@@
We first define a grid of regularization values, $\Lambda$, from which the optimal $\lambda$ for Lasso is to be identified. For this purpose, we use as the default choice the $\lambda$ values generated by the cv.glmnet function from the glmnet package in R (Friedman et al., 2010), applying a 10-fold cross-validation to the complete dataset $\mathcal{D}$. Alternative regularization values can also be employed. However, a comparison of different methods for generating the regularization grid falls outside the scope of this paper. The goal is to identify the optimal regularization value $\lambda_{\text{stable}}\in\Lambda$ that yields highly stable outcomes with the least possible loss in terms of predictive ability.
The following procedure is applied to each regularization value $\lambda\in\Lambda$. In each iteration of the stability selection process, a sub-sample, comprising half the size of the original dataset $\mathcal{D}$, is randomly selected. The Lasso model is then fitted to this sub-sample using the fixed value of $\lambda$, and the binary selection outcomes are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the number of sub-samples. Therefore, in the end, we obtain $|\Lambda|$ distinct selection matrices $M(\lambda)$. In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is identified as part of the signal set $\hat{S}(\lambda)$ when Lasso is applied to the $b$th sub-sample given $\lambda$. In contrast, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is classified as belonging to the noise set $\hat{N}(\lambda)$ after applying Lasso to the $b$th sub-sample given $\lambda$.
@@
$\hat{\Phi}(M(\lambda)):=1-\frac{\frac{1}{p}\sum_{j=1}^{p}s_{j}^{2}}{\frac{q(\lambda)}{p}\left(1-\frac{q(\lambda)}{p}\right)};\quad\lambda\in\Lambda,$ (2)
where $s_{j}^{2}$ denotes the unbiased sample variance of the binary selection statuses of the $j$th variable, while $q(\lambda)$ denotes the average number of variables selected under the regularization parameter $\lambda$. The stability estimator $\hat{\Phi}(\cdot)$ is bounded by $\left[-\frac{1}{B-1},1\right]$ (Nogueira et al., 2018); the larger the $\hat{\Phi}(\cdot)$, the more stable is $M(\lambda)$. To interpret the stability values of Equation (2), Nogueira et al. (2018) adopted the guidelines proposed by Fleiss et al. (2004). According to Nogueira et al. (2018), stability values that exceed $0.75$ indicate excellent agreement between selections beyond what would be expected by random chance, while values below $0.4$ signify poor agreement between them. Based on Nogueira et al. (2018), stability values that reside between these thresholds are categorized as indicative of intermediate to good stability.
@@
$\lambda_{\text{stable}}:=\min\left\{\lambda\in\Lambda\mid\hat{\Phi}(M(\lambda))\geq 0.75\right\}.$ (3)
The threshold value of $0.75$ is somewhat arbitrary and may not be attainable in certain situations. As we will see in Section 4, $\lambda_{\text{stable}}$ may not exist in certain practical applications due to procedural instability, which prevents the stability values from exceeding $0.75$. In such cases, we propose
$\lambda_{\text{stable-1sd}}:=\min\left\{\lambda\in\Lambda\mid\hat{\Phi}(M(\lambda))\geq\max_{\lambda\in\Lambda}\hat{\Phi}(M(\lambda))-\text{SD}_{\lambda\in\Lambda}\left(\hat{\Phi}(M(\lambda))\right)\right\},$ (4)
where SD()($\cdot$)( ⋅ ) represents the standard deviation function.
Both upper-bounds of the Per-Family Error Rate, mentioned in Section 1, provided by Meinshausen and Bühlmann (2010) and Shah and Samworth (2013), depend on the values of $\lambda$ and $\pi_{\text{thr}}$. The calibration of these two values requires the arbitrary selection of one of these two parameters, which can be challenging to justify (Bodinier et al., 2023). By adopting the optimal regularization value, that is $\lambda=\lambda_{\text{stable}}$ or $\lambda=\lambda_{\text{stable-1sd}}$, the upper-bound can be explicitly tailored to the data owner’s preferences. If a pre-determined $\pi_{\text{thr}}$ is chosen, the Per-Family Error Rate is deterministically obtained; alternatively, if a fixed pre-determined Per-Family Error Rate value is desired, the corresponding $\pi_{\text{thr}}$ is enforced.
@@
where $q_{\Lambda}$ represents the average number of selected variables over the regularization grid $\Lambda$. In a manner consistent with the approach of Bodinier et al. (2023), we can derive a point-wise control version of Equation (5) by restricting $\Lambda$ to a single value, $\lambda$, without affecting the validity of the original equation. Consequently, we can replace $q_{\Lambda}$ with $q(\lambda)$, as introduced in Equation (2). By employing $\lambda_{\text{stable}}$ within this formula and utilizing $q(\lambda_{\text{stable}})$, we establish a two-way control: fixing $\pi_{\text{thr}}$ allows us to determine the upper-bound, and vice versa. A similar approach can be applied to the upper-bounds proposed by Shah and Samworth (2013).
@@
$\hat{S}^{\text{stable}}:=\left\{j\mid\hat{\Pi}_{j}^{\lambda_{\text{stable}}}\geq\pi_{\text{thr}}\right\};\quad j=1,\ldots p.$ (6)
Equation (6) represents the variables with selection frequencies higher than $\pi_{\text{thr}}$ under the $\lambda_{\text{stable}}$, which corresponds to the highly stable outcomes. We term this selection criterion ‘Stable Stability Selection’ because it leverages highly stable results to facilitate variable selection. If $\lambda_{\text{stable}}$ does not exist, $\hat{S}^{\text{stable-1sd}}$ can be defined similarly.
As outlined in Section 1, as the number of sub-samples increases, the convergence of $\hat{\Phi}(\cdot)$ to the true stability $\Phi$ is ensured. To illustrate this, we define $M^{(t)}(\lambda)$ as the matrix containing the selection results from the first $t$ sub-samples for a given $\lambda\in\Lambda$, that is, the first $t$ rows of $M(\lambda)$. The objective is to evaluate the stability of $M^{(t)}(\lambda)$ across successive sub-samples in order to determine the appropriate cut-off point for the process, that is, the number of sub-samples required to achieve convergence in stability. Beyond this threshold, additional sub-samples do not significantly change the stability of $M(\lambda)$. We propose plotting the stability values of the selection matrix over the sequential sub-sampling, that is $\hat{\Phi}(M^{(t)}(\lambda));\;t=2,3,\ldots B$, against $t$ to monitor the convergence status of the stability values. Given that the asymptotic distribution of $\hat{\Phi}(\cdot)$ is established, a confidence bound can be drawn along the curve to reflect the inherent uncertainty of the stability estimator.